Some oxygen is collected in a 0.250 L container at 25.0C and 720.mmHg. How many moles of oxygen is

this?


1
Expert's answer
2021-05-19T02:26:39-0400

We know that

1atm = 760 mmHg

So 720 mmHg=0.94atm

T=25+273=298K

R= 0.082L·atm/K·mol

V=0.025L

We will use

"PV=nRT"

0.94atm×0.25L=n×0.082L·atm/K·mol×298K

n=0.0096 mole